OpenStudy (anonymous):

Let f (n) denote the nth term of a sequence of integers given by the equation f(n) = f(n-1) +f(n-2) for n > 2 and f(1) = 1 and f(2) = 1, then using principle of mathematical induction, show that √ 5 f( n ) = {(1+ √ 5) / 2} n -- {(1 -- √ 5) / 2}n for all n > = 1

OpenStudy (anonymous):

\[f(n)=f(n-1)+f(n-2),~~n>2\\ f(1)=1, ~f(2)=1\] And you have to show that \[\sqrt5f(n)=\frac{1+\sqrt5}{2}n-\frac{1-\sqrt5}{2}n,~~\text{for }n\ge1\] (You have \(\left\{f(n)\right\}=\left\{1,1,2,3,5,8,13,21,...\right\}\), which happens to be the Fibonacci sequence.) First show that the equality holds for \(n=1\): \[\sqrt5f(1)=\frac{1+\sqrt5}{2}-\frac{1-\sqrt5}{2}\\ \sqrt5=\frac{\sqrt5+\sqrt5}{2}\\ \sqrt5=\sqrt5\] Now assume it holds for \(n=k\): \[\sqrt5f(k)=\frac{1+\sqrt5}{2}k-\frac{1-\sqrt5}{2}k\] and, under this assumption, show that it holds for \(n=k+1\): \[\sqrt5f(k+1)=\frac{1+\sqrt5}{2}(k+1)-\frac{1-\sqrt5}{2}(k+1)\] What do you know about \(f(k+1)\), and how can you rewrite the last equation so that it uses the hypothesis that the equality holds for \(k\) ?